Twenty meters of wire is available for fencing off a flower bed in the form of a circular sector. Then the maximum area of the flower bed is

Grade:12

1 Answers

Arun
5 years ago

Perimeter of a sector is P = (2 + theta)R, and area is A = ½·R²·theta

theta = P/R -2, substituting in area equation

A = P·R/2 - r², we wish to maximize area

dA/dR = P/2 - 2·R —-> R = P/4 = 5 m, and theta = 2 rad

A = 25 m²